OpenStudy (anonymous):

negative twelve minus negative nineteen equals blank negative thirty-one negative seven 7 thirty-one

OpenStudy (shamil98):

-12 - (-19 ) = -12 + 19 = ? - x - = + that's why you add 19

OpenStudy (anonymous):

-7 right

OpenStudy (shamil98):

no. -12 + 19 is the same as 19 - 12 ..

OpenStudy (anonymous):

-31

OpenStudy (anonymous):

or is it 7

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

19 - 12 = 7 ...